actory Overhead Rates, Entries, and Account Balance Sundance Solar Company operates two factories. The company applies factory overhead to jobs on the basis of machine hours in Factory 1 and on the basis of direct labor hours in Factory 2. Estimated factory overhead costs, direct labor hours, and machine hours are as follows: Factory 1 Factory 2 Estimated factory overhead cost for fiscal year beginning March 1 $460,460 $624,000 Estimated direct labor hours for year 10,400 Estimated machine hours for year 20,930 Actual factory overhead costs for March $36,760 $54,140 Actual direct labor hours for March 940 Actual machine hours for March 1,630 a. Determine the factory overhead rate for Factory 1. $ per machine hour b. Determine the factory overhead rate for Factory 2. $ per direct labor hour c. Journalize the entries to apply factory overhead to production in each factory for March. Factory 1 Factory 2 d. Determine the balances of the factory overhead accounts for each factory as of March 31, and indicate whether the amounts represent overapplied factory overhead or underapplied factory overhead. Factory 1 $ Factory 2 $
Answer #1

a) Factory overhead rate for Factory 1 = 460460/20930 = 22 per machine hour

b) Factory overhead rate for Factory 2 = 624000/10400 = 60 per labor hour

c) Journal entry-Factory 1

Date account and explanation Debit Credit
Work in process (1630*22) 35860
Factory overhead 35860
(To record applied overhead

c) Journal entry-Factory 2

Date account and explanation Debit Credit
Work in process (940*60) 56400
Factory overhead 56400
(To record applied overhead

d) Factory 1 = 36760-35860 = 900 Under applied

Factory 2 = 54140-56400 = 2260 Over applied